# Aligning Large Language Models with Self-generated Preference Data

### v1

@@ 6 Conclusion
In this paper, we proposed Selfee, a method that can efficiently improve the alignment of LLMs using minimal human-labeled preference data. Our main contributions include the development of an effective data expansion method with the direct preference judgment method and preference learning algorithm with the self-refinement of (potentially) noise preference. We demonstrate the effectiveness of Selfee by fine-tuning the recent LLMs with the various setups, and observing the significant improvements when evaluating them on the commonly used benchmarks.
Limitation and future work. In the experiments, Selfee has shown the tendency to increase the responses’ length. We demonstrated that the improvement by Selfee is not a simple result of such length increase, by observing the increase of win rate under a length length-controlled setup or MT-bench. However, depending on the user, this behavior could be dispreferred. In this sense, focusing on mitigating this bias during the self-improving alignment will be an interesting future direction, and can enhance the robustness and generalizability of Selfee across more diverse scenarios.
Appendix A Societal Impact
Selfee enables efficient human preference learning, allowing for cost-effective training of models in data-scarce or domain-specific areas. Our framework supports alignment learning in various fields, including multilingual language learning and preferences beyond human helpfulness. Consequently, it could contribute to facilitating the widespread adoption of LLM technology across diverse sectors. By lowering the barriers to alignment learning, Selfee makes it more accessible to a broader audience. However, the widespread availability of this technology also brings potential risks. The reduced cost of training models could enable malicious actors to misuse the technology, leading to societal issues. Therefore, it is crucial to implement ethical considerations and safety measures when deploying Selfee technology to mitigate these risks.
@@ B.1 SFT model setup
Mistral. For supervised fine-tuning, Ultrachat dataset [8] is used [^9] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k, batch size was set 128, total epoch was 1, and the learning rate was $2\times 10^{-5}$. It employed Adam optimizer [17] and a cosine learning rate scheduler with a warm-up phase corresponding to 10% of the total training steps.
@@ B.2 Baselines experiment setup
Zephr-7b-$\beta$. Zephr-7b-$\beta$ [40], which is compared in Table 1, was trained using the same pre-trained model (mistral-7b-0.1v [15]) and the same SFT data (Ultrachat [8]), but there are marginal differences compared to our setup. Specifically, Zephr-7b-$\beta$ used the batch size of 512, but 128 was used for the model in our experiments. As a result, the SFT model of Zephr-7b-$\beta$ [^10] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta, according to AlpacaEval2.0, showed a length control win rate of 8.14% and a win rate of 5.26%, approximately 0.5% better than the SFT model used in our experiments (8.14% vs 7.58%, 5.26% vs 4.72%). In addition, regarding the preference dataset, Zephr-7b-$\beta$ was trained using 100% of the original Ultrafeedback [7]  [^11] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized but we use 3.3% of its cleaned version [^12] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned. Consequently, the results under the two datasets might be slightly different, and the impact of these method differences has been specifically denoted in [14]; here, Ultrafeedback clean shows approximately a 0.9% improvement in performance based on the AlpacaEval2.0 win rate (from 11.33% to 12.2%). Consequently, although Zephr-7b-$\beta$ is not a perfectly strict comparison, we believe the minor variations are relatively small compared to the differences between Selfee and Zephr-7b-$\beta$, and hence the comparison results are meaningful.
@@
LLM-as-Judgement. For LLM-as-judge, we used an SFT model to employ Consitual AI’s pairwise comparison prompt for judging preferences [2]. Preference is measured by comparing the logprob value of the token output as input to the following prompt (Listing 7). To ensure fair comparison and prevent low judgment performance, evaluation instructions were created using seed preference data which is the same form as Consitual AI’s pairwise comparison. (Listing 7) Using these, additional SFT learning is performed to obtain an independent LLM-as-judge model. For this supervised fine-tuning, we set the batch size 32, total epoch is 3, and the learning rate was $2\times 10^{-5}$. We employed Adam optimizer and a cosine learning rate scheduler with a warm-up phase corresponding to 10% of the total training steps.
@@
Reward model judgment. For the reward model baseline, we selected PairRM [16] due to its high performance on AlpacaEval2.0 [35, 45]. Unlike SELFEE, which was trained on only 2K gold label data, PairRM was trained on a large-scale dataset. The training data for PairRM includes the following:
@@
The total number of pairwise samples in this training data is approximately 500K, compared to 2K for SELFEE. Specifically, the summarize_from_feedback dataset contributes 179K samples, and the hh-rlhf dataset contributes 161K samples, making up a significant portion of the total.
B.3 Implementation details
@@
Table 7: MT-Bench. Evaluation results on MT-bench with different models. Selfeeinst and Selfeephi denotes the fine-tuned Mistral-7B-Instruct-v0.1 and Phi-2 SFT with Selfee, respectively.
Models Gold Label (%) Avg Score Mistral-7B-Instruct-v0.1 - 6.78 Selfeeinst (Ours) - 7.12 Phi-2 SFT - 5.35 Phi-2 DPO 3.3 6.16 Selfeephi (Ours) 3.3 6.33
@@
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Consider the following conversation between a human and an assistant: [prompt] Please choose the response that is the most helpful, honest, and harmless. Options: (A) [response 1] (B) [response 2] The answer is: ”’ Prompt used for judging preferences. {listing}[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Consider the following conversation between a human and an assistant: [Horses are hybrids of which two animals?] Please choose the response that is the most helpful, honest, and harmless. Options: (A) [Horses, scientifically known as Equus caballus, are not the result of hybridization between two other animals. They are considered a species of their own within the Equidae family, which also includes other equines such as donkeys, mules, and zebras. Horses are believed to have evolved from a small, multi-toed forest-dwelling animal known as Eohippus, which lived around 60 million years ago. Over millions of years, these horses evolved into the large, single-toed animals we know today.] (B) [Horses are not hybrids of any other animals. They are a unique species and are not a product of cross-breeding between different animals. Horses are domesticated animals and have been bred and developed over thousands of years by humans for various purposes such as transportation, riding, and agriculture.] The answer is: (A)
@@ Appendix C More Evaluation Results in MT-bench
In this section, we further provide the additional evaluation results on MT-bench [16], which couldn’t be presented in the main text, due to the space issue.
@@
Figure 5: More evaluation results with MT-bench.
@@
We first present (a) task-wise evaluation results and (b) iteration-wise average improvement in Figure 5. As shown in Figure 5(a), Selfee consistently improves the performance in various tasks. Notably, there is almost no gain in Coding and degradation in Math. We remark that this phenomenon is commonly observed in the relevant literature [22], which indicates that different training [41] or inference [44] schemes might be necessary to improve the performance in these tasks.
@@
Next, in Figure 5(b), one can observe that the average performance on the MT-bench is increased with more iterations. Specifically, while the iterative DPO using PairRM shows the best performance until iteration 2, Selfee outperforms it in iteration 3. It demonstrates the effectiveness of our framework for iteratively improving the alignment of LLM.
@@
Lastly, we present additional evaluation results on MT-bench, not presented in Table 6. Specifically, we measure the performances of Phi-2 variants and Mistral-7B-Instruct-v0.1 variants, which presented in Table 4 and Figure 4. The results are presented in Table 7. As one can see, Selfee consistently yields the improvement across different backbones of Mistral-7B-Instruct-v0.1 and Phi-2.
Appendix D Qualitative Examples
@@
In this section, we present examples of the responses from the models trained under different methods (SFT, DPO, and Selfee), for a qualitative comparison. The results are presented in Figures 6, 7, and 8. As one can see in the examples, the model trained under Selfee provides more detailed and informative responses to various questions.
@@
Figure 6: Response in AlpacaEval2.0
@@
Figure 7: Response in AlpacaEval2.0
@@
Figure 8: Response in AlpacaEval2.0

### v2

@@ 6 Conclusion
In this paper, we proposed SPA, a method that can efficiently improve the alignment of LLMs using minimal human-labeled preference data. Our main contributions include the development of an effective data expansion method with the direct preference judgment method and a preference learning algorithm with the self-refinement of (potentially) noise preference. We demonstrate the effectiveness of SPA by fine-tuning the recent LLMs with the various setups, and observing the significant improvements when evaluating them on the commonly used benchmarks, AlpacaEval 2.0 and MT-Bench. We expect SPA to make significant contributions to future research and practical applications, especially when the human-labeled preference is hard to collect. Limitations and societal impacts are further discussed in Appendix A.
Reproducibility Statement
For the reproducibility of our results, we have provided a detailed description of our methods and experimental setups in Section 5.1 and Appendix B. We also confirmed the robustness of our results through the experiment (Table 4). In addition, to further facilitate the reproduction, we will release our codes and the checkpoints for the trained models.
Acknowledgments
This work was mainly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2021-0-02068, Artificial Intelligence Innovation Hub, 50%; RS-2022-II220959, Few-shot Learning of Casual Inference in Vision and Language for Decision Making, 50%).
Appendix A Limitation and Societal Impact
A.1 Limitation and future work
In the experiments, SPA has shown the tendency to increase the responses’ length (please see Appendix D for the relevant results and discussions). We demonstrated that the improvement by SPA is not a simple result of such length increase, by observing the increase of win rate under a length length-controlled setup or MT-bench. However, depending on the user, this behavior could be dispreferred. In this sense, focusing on mitigating this bias during the self-improving alignment will be an interesting future direction, and can enhance the robustness and generalizability of SPA across more diverse scenarios.
A.2 Societal impact
SPA enables efficient human preference learning, allowing for cost-effective training of models in data-scarce or domain-specific areas. Our framework supports alignment learning in various fields, including multilingual language learning and preferences beyond human helpfulness. Consequently, it could contribute to facilitating the widespread adoption of LLM technology across diverse sectors. By lowering the barriers to alignment learning, SPA makes it more accessible to a broader audience. However, the widespread availability of this technology also brings potential risks. The reduced cost of training models could enable malicious actors to misuse the technology, leading to societal issues. Therefore, it is crucial to implement ethical considerations and safety measures when deploying SPA technology to mitigate these risks.
@@ B.1 SFT model setup
Mistral. For supervised fine-tuning, Ultrachat dataset (Ding et al., 2023) is used [^11] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k, batch size was set 128, total epoch was 1, and the learning rate was $2\times 10^{-5}$. It employed Adam optimizer (Kingma & Ba, 2015) and a cosine learning rate scheduler with a warm-up phase corresponding to 10% of the total training steps.
@@
LLaMA-3 and Phi-3. As described in Section 5.3, we use the generically instruct-tuned versions for both LLaMA-3-8B and Phi-3-14B, as there are no SFT models tuned on Ultrachat dataset.
@@ B.2 Baselines experiment setup
Zephyr-7b-$\beta$. We implemented Zephyr-7b-$\beta$ (Tunstall et al., 2023), which is compared in Table 1, according to recipes. Our Zephyr-7b-$\beta$ was trained using the same pre-trained model (mistral-7b-0.1v (Jiang et al., 2023a)) and the same SFT data (Ultrachat (Ding et al., 2023)), but there are marginal differences compared with recipes. We use SFT  [^12] https://huggingface.co/alignment-handbook/zephyr-7b-sft-full models which trained with different recipes. Specifically, Zephyr-7b-$\beta$’s SFT used the batch size of 512, but 128 was used for the ours SFT model. In addition, regarding the preference dataset, Zephyr-7b-$\beta$ was trained using the original Ultrafeedback (Cui et al., 2023)  [^13] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized but we use cleaned version [^14] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned. These changes in training data and the SFT model were aligned with SPA to ensure a fair comparison.
@@
LLM-as-Judgement. For LLM-as-judge, we used an SFT model to employ Consitual AI’s pairwise comparison prompt for judging preferences (Bai et al., 2022a). Preference is measured by comparing the logprob value of the token output as input to the following prompt (Listing 8). To ensure fair comparison and prevent low judgment performance, evaluation instructions were created using seed preference data which is the same form as Consitual AI’s pairwise comparison. (Listing 8) Using these, additional SFT learning is performed to obtain an independent LLM-as-judge model. For this supervised fine-tuning, we set the batch size 32, total epoch is 3, and the learning rate was $2\times 10^{-5}$. We employed Adam optimizer and a cosine learning rate scheduler with a warm-up phase corresponding to 10% of the total training steps.
@@
Reward model judgment. For the reward model baseline, we selected PairRM (Jiang et al., 2023b) due to its high performance on AlpacaEval 2.0 (Snorkel, 2024; Wu et al., 2024). Unlike SPA, which was trained on only 2K gold label data, PairRM was trained on a large-scale dataset. The training data for PairRM includes the following:
@@
The total number of pairwise samples in this training data is approximately 500K, compared to 2K for SPA. Specifically, the summarize_from_feedback dataset contributes 179K samples, and the hh-rlhf dataset contributes 161K samples, making up a significant portion of the total.
B.3 Adjusted experimental setups for different LLMs
In Table 5, we conduct the experiments with different LLMs. As they exhibit different characteristics from the difference in backbone and sizes, we slightly adjusted the experimental setups while keeping most identical to the setups in Section 5.1.
Phi-2. We slightly adjust the learning rate to accommodate the different characteristics of the Phi-2 ($5\times 10^{-6}$). In addition, due to the smaller size of the Phi-2, we observe that performance improvements were not evident beyond iteration 2. Therefore, we present the results of iteration 1.
LLaMA-3 and Phi-3. We slightly adjust the learning rate to accommodate the different characteristics of models ($1\times 10^{-5}$). We conduct 1 epoch for the initial DPO training and maintain $\beta=0.01$ throughout the entire training process. Since performance improvement has been only observed up to iteration 2 in Section 5.2. we conduct the experiments up to iteration 2 for these models.
B.4 Implementation details
@@
Table 8: MT-Bench. Evaluation results on MT-bench with different models. SPAinst and SPAphi denotes the fine-tuned Mistral-7B-Instruct-v0.1 and Phi-2 SFT with SPA, respectively.
Models Gold Label (%) Avg Score Mistral-7B-Instruct-v0.1 - 6.78 SPAinst (Ours) - 7.12 Phi-2 SFT - 5.35 Phi-2 DPO 3.3 6.16 SPAphi (Ours) 3.3 6.33
@@
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Consider the following conversation between a human and an assistant: [prompt] Please choose the response that is the most helpful, honest, and harmless. Options: (A) [response 1] (B) [response 2] The answer is: ”’ Prompt used for judging preferences.
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Consider the following conversation between a human and an assistant: [Horses are hybrids of which two animals?] Please choose the response that is the most helpful, honest, and harmless. Options: (A) [Horses, scientifically known as Equus caballus, are not the result of hybridization between two other animals. They are considered a species of their own within the Equidae family, which also includes other equines such as donkeys, mules, and zebras. Horses are believed to have evolved from a small, multi-toed forest-dwelling animal known as Eohippus, which lived around 60 million years ago. Over millions of years, these horses evolved into the large, single-toed animals we know today.] (B) [Horses are not hybrids of any other animals. They are a unique species and are not a product of cross-breeding between different animals. Horses are domesticated animals and have been bred and developed over thousands of years by humans for various purposes such as transportation, riding, and agriculture.] The answer is: (A)
@@ Appendix C More Evaluation Results in MT-Bench
In this section, we further provide the additional evaluation results on MT-Bench (Zheng et al., 2023), which couldn’t be presented in the main text, due to the space issue.
@@
Figure 5: MT-bench Evaluation. More evaluation results with MT-bench.
@@
We first present (a) task-wise evaluation results and (b) iteration-wise average improvement in Figure 5. As shown in Figure 5(a), SPA consistently improves the performance in various tasks. Notably, there is almost no gain in Coding and degradation in Math. We remark that this phenomenon is commonly observed in the relevant literature (Lin et al., 2024), which indicates that different training (Wang et al., 2023a) or inference (Wei et al., 2022b) schemes might be necessary to improve the performance in these tasks.
@@
Next, in Figure 5(b), one can observe that the average performance on the MT-bench is increased with more iterations. Specifically, while the Iterative DPO using PairRM shows the best performance until iteration 2, SPA (without DND) outperforms it in iteration 3. It demonstrates the effectiveness of our framework for iteratively improving the alignment of LLM.
Table 9: Ablation study including MT-Bench. Evaluation results on AlpacaEval 2.0 and MT-Bench with iteratively trained models (from SFT) under different methodological configurations of SPA. DE, SR, DND are abbreviations of data expansion, self-refinement, and de-coupled noise detection, respectively. The best scores are highlighted with bold.
AlpacaEval 2.0 MT-Bench Methods DE SR DND Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) Avg. Score (0-10) SFT - - - 7.58 4.72 6.34 DPO - - - 9.03 7.68 6.81 14.41 19.91 6.86 SPA (Ours) 14.7 19.94 7.09 15.39 21.13 6.94
@@
In addition, we measure the performances of Phi-2 variants and Mistral-7B-Instruct-v0.1 variants on MT-Bench in Table 8; these models are presented in Table 5 and Figure 4, respectively. As one can see, SPA consistently yields the improvement across different backbones of Mistral-7B-Instruct-v0.1 and Phi-2. Lastly, we present the full results of the ablation study (presented in Table 6) that includes the evaluation results on MT-Bench, in Table 9.
Appendix D More Quantitative Results
In this section, we present more quantitative results to demonstrate the effectiveness of SPA.
Table 10: SPA with length regularization. Evaluation results on AlpacaEval2.0 with different variants of Mistral-7B-v0.1 from SPA and the additional length regularization term.
Models Gold Label (%) Len-control. Win Rate Win Rate vs. GPT-4 Avg. len (# chars) Mistral-7B-v0.1 - 0.17 0.50 5692 SFT - 7.58 4.72 901 DPO 3.3 9.03 7.68 1802 Zephyr-7b-$\beta$ 100 11.75 10.03 1552 SPA (Original, Iter. 1) 3.3 11.88 12.95 2150 SPA (Modified, Iter. 1) 3.3 11.39 12.31 2013 SPA (Original, Iter. 2) 3.3 16.23 19.94 2749 SPA (Modified, Iter. 2) 3.3 14.46 18.23 2448
Mitigating length bias with SPA. Here, we provide a discussion of the relevant experimental results about the length bias present in SPA. During the experiments, we observe that LLMs trained with SPA tend to generate longer responses (see 10), which could be dispreferred depending on the user. Regarding this, we first emphasize that the improvement with SPA is not merely due to longer outputs, as shown by the significant gains in the length-controlled win rate in all experiments in Section 5.
Nevertheless, to further address the concerns regarding this issue, we further investigate whether previously researched length control techniques can be easily integrated into SPA. Specifically, we apply the length penalty approach from RLHFlow Dong et al. (2024). This method heuristically reduces the reward model’s reward based on the output length (Eq. 13) during preference labeling. We utilze hyperparamter $\alpha$ between $0.001$ to $0.01$ that minimize the length increase. The results, shown in Table 10, indicate that this modification successfully reduces the average length while largely preserving the performance improvements from SPA. These results demonstrate that SPA can be easily integrated with existing research related to length control.
$r_{\tt penalty}(x,y)=r(x,y)-\alpha|y|$ (13)
Table 11: LLM-as-Judgment with model from previous iteration. Evaluation results on AlpacaEval 2.0 with different variants of Mistral-7B-v0.1.
Methods Len-control. Win Rate (%) Win Rate vs. GPT-4 (%) LLM-as-judge (Iter. 1) 8.88 8.01 LLM-as-judge (Iter. 2, orig) 9.49 8.46 LLM-as-judge (Iter. 2, prev. init) 9.74 10.09 SPA (Iter. 2, ours) 15.46 19.91
LLM-as-Judge with the model from previous iteration. For the LLM-as-Judgement baseline, we used a fixed model fine-tuned specifically for evaluating preferences between responses, as described in Appendix B.2. On the other hand, at each iteration, the proposed SPA utilizes the previous iteration’s model, $\pi_{i-1}$, for the direct preference judgment. Therefore, one can be curious how this difference affect to the performance.
To answer this, we conducted new experiments; at the 2nd iteration, the evaluation model was initialized with the resulting model from the 1st iteration and fine-tuned with the constructed preference evaluation dataset using the seed preference data (see details in Appendix B.2). The evaluation results (denoted as LLM-as-Judge (Iter. 2, prev. init)) on AlpacaEval 2.0 are presented in Table 11, along with other methods at the 2nd iteration. While this approach yielded improved alignment compared to the fixed model, SPA still significantly outperformed this baseline. This underscores that the superior performance of SPA arises from its novel preference evaluation techniques rather than the specific evaluation model used.
Appendix E Qualitative Examples
@@
In this section, we present examples of the responses from the models trained under different methods (SFT, DPO, and SPA), for a qualitative comparison. The results are presented in Figures 6, 7, and 8. As one can see in the examples, the model trained under SPA provides more detailed and informative responses to various questions.
@@
Figure 6: Example of responses in AlpacaEval 2.0.
@@
Figure 7: Example of responses in AlpacaEval 2.0.
@@
Figure 8: Example of responses in AlpacaEval 2.0.